OpenStudy (anonymous):

Write a definite integral which can be approximated to be the Riemann Sum of: 1/50 ( sqrt of (1/50) + sqrt of (2/50) + sqrt of (3/50)....sqrt of (50/50) ).

OpenStudy (anonymous):

ill write it with teh equation thing also

OpenStudy (anonymous):

\[1/50* (\sqrt{1/50} + \sqrt{2/50}+ \sqrt{3/50}...\sqrt{50/50}) \]

OpenStudy (zarkon):

you have any idea of where to start?

OpenStudy (zarkon):

n=10?

OpenStudy (zarkon):

looks like you have 50 to me

OpenStudy (zarkon):

\[\sum_{i=1}^{50}\sqrt{\frac{i}{50}}\frac{1}{50}\]

OpenStudy (anonymous):

so the function is sqrt of 1/50 times 1/50?

OpenStudy (zarkon):

your function is \[f(x)=\sqrt{x}\]

OpenStudy (zarkon):

\[\Delta x=\frac{1}{50}\] \[x_i^*=a+i\times\Delta x=\frac{i}{50}\] then you can write the above as \[\sum_{i=1}^{50}f(x_i^*)\Delta x\]

OpenStudy (zarkon):

\[\int\limits_{a}^{b}f(x)dx=\lim_{n\to\infty}\sum_{i=1}^{n}f(x_i^*)\Delta x\]

OpenStudy (anonymous):

so wat was a in this case? 1?

OpenStudy (zarkon):

a=0

OpenStudy (zarkon):

b=1

OpenStudy (zarkon):

\[\Delta x=\frac{b-a}{n}=\frac{1-0}{50}=\frac{1}{50}\]

OpenStudy (anonymous):

srry i cant read that clearlythe limit says as wat approaches infinity?

OpenStudy (zarkon):

yes...as n goes to infinity then we get the definite integral

OpenStudy (anonymous):

this makes sense now! thanks for explaining teh steps b.c i was confused with how u got the asnwer

OpenStudy (anonymous):

srry earler for sum reason i was talkin about another problem which is why i said n = 10 n etc..

OpenStudy (zarkon):

the integral you will get in the end is \[\int\limits_{0}^1\sqrt{x}dx\]

OpenStudy (anonymous):

why is taht?

OpenStudy (zarkon):

look from where I said "your function is \[f(x)=\sqrt{x}\] "

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

\[1/50* (\sqrt{1/50} + \sqrt{2/50}+ \sqrt{3/50}...\sqrt{50/50})\approx.676095\] \[\int\limits_{0}^1\sqrt{x}dx=.\overline{6}\]

OpenStudy (anonymous):

thats really cool

OpenStudy (anonymous):

calc is mucch more fun n interesting when i actually understand it

OpenStudy (zarkon):

yep...it's good stuff

OpenStudy (anonymous):

i tend to b kinda slow especially wen things arent done step by step...srry

OpenStudy (zarkon):

np